OpenStudy (anonymous):

A K2CrO4 solution was standardized by titrating 0.285g of Cu, 29.60 mL were used. Calculate the normality of the solution. (CrO4)2- --> (Cr)3+ Cu --> (Cu)2+

OpenStudy (anonymous):

do u know how to calculate normality of a solution? let me know at which step u r confused

OpenStudy (anonymous):

I have a vague idea of some of the equations, but I don't know how to put it together. I am confused about the equations below it and their relevance.

OpenStudy (cuanchi):

You have a REDOX titration here, you have to find how many electrons are transferred between the elements (compounds) in the reaction. Normality is defined in this case has the equivalents/L of solution. Eq/L One equivalent is going to be equal to the molecular mass divided by the number of electrons transferred. 1) write the redox reaction, balance it by the half reaction method and calculate how many e- has been transferred. (if you need help let me know) 2) calculate how many moles are in 0.285g of Cu. n= m/MM 3) calculate how many Eq are in the "n" mol of Cu. Eq=n/number of e- 4) because the reactions are Eq Cu to Eq (CrO4)2- the same number of Eq of Cu you will have of Eq of (CrO4)2- 5) convert the volume of solution from mL to L and calculate the Normality, N= Eq (CrO4)2-/volume of (CrO4)2- in liters (N=Eq/V(L))

OpenStudy (cuanchi):

Normality is an ambiguous measure of the concentration of a solution. It needs a definition of the equivalence factor, which depends on the definition of equivalents. The same solution can possess different normalities for different reactions. The definition of the equivalence factor varies depending on the type of chemical reaction that is discussed: It may refer to equations, bases, redox species, precipitating ions, or isotopes. For example, a solution of MgCl2 that is 2 N with respect to a Cl− ion, is only 1 N with respect to an Mg2+ ion. Since f_\mathrm{eq} may not be unequivocal, IUPAC and NIST discourage the use of normality. https://en.wikipedia.org/wiki/Equivalent_concentration#Criticism